# Artificial Inductive Bias for Synthetic Tabular Data Generation in Data-Scarce Scenarios

### v1

@@ Abstract
While synthetic tabular data generation using Deep Generative Models (DGMs) offers a compelling solution to data scarcity and privacy concerns, their effectiveness relies on substantial training data, often unavailable in real-world applications. This paper addresses this challenge by proposing a novel methodology for generating realistic and reliable synthetic tabular data with DGMs in limited real-data environments. Our approach proposes several ways to generate an artificial inductive bias in a DGM through transfer learning and meta-learning techniques. We explore and compare four different methods within this framework, demonstrating that transfer learning strategies like pre-training and model averaging outperform meta-learning approaches, like Model-Agnostic Meta-Learning, and Domain Randomized Search. We validate our approach using two state-of-the-art DGMs, namely, a Variational Autoencoder and a Generative Adversarial Network, to show that our artificial inductive bias fuels superior synthetic data quality, as measured by Jensen-Shannon divergence, achieving relative gains of up to 50% when using our proposed approach. This methodology has broad applicability in various DGMs and machine learning tasks, particularly in areas like healthcare and finance, where data scarcity is often a critical issue.
@@ 1 Introduction
Recent advances in deep learning have led to the development of powerful Deep Generative Models (DGMs). These models excel at learning and representing complex, high-dimensional distributions, allowing them to sample new data points realistically. This capability has driven remarkable progress in various domains, including image generation [1], text generation [2], and video generation [3].
Tabular data has emerged as a data format of increasing interest within DGMs. Structured in a familiar spreadsheet-like format with rows and columns, tabular data are a fundamental cornerstone for information storage and analysis across diverse fields. Researchers are actively investigating techniques for generating realistic and informative tabular data, as evidenced by the growing body of work in this area [4, 5, 6]. Generative Adversarial Networks (GANs) are currently the most prominent architecture for tabular data generation. Within this domain, CTGAN [7] is a widely recognized model. However, the exploration of alternative methodologies is ongoing, as evidenced by surveys [8] that highlight diverse approaches [9, 10, 11]. Beyond GANs, there are promising approaches including Variational Autoencoders (VAEs) [12], methods based on diffusion models [13], and techniques that leverage language models [14], to mention some. Focusing on VAEs for tabular data generation, TVAE [7] stands as a foundational model. TVAE encodes tabular data into a latent space, a lower-dimensional representation that captures the essential features of the data. By sampling from this latent space and subsequently decoding the samples, TVAE facilitates the generation of new data points in the original tabular format. A recent advancement [15] builds upon TVAE by incorporating a Bayesian Gaussian Mixture model (BGM) within the VAE architecture. This novel approach outperforms the state-of-the-art models, including CTGAN and TVAE, on different validation metrics. This superior performance comes from the BGM integration enabling high-quality sampling from the VAE’s latent space. This addresses a limitation of TVAE, which assumes a Gaussian latent space for generation, which may not happen in practice [16]. As real-world data often deviates from this distribution, the BGM integration leads to improved performance.
@@
The importance of sufficient data for DGM training cannot be overstated. Studies using popular DGMs, such as the CTGAN, use datasets ranging from 10k to 481k training instances. This contrasts starkly with the practices observed in numerous domains. For example, well-established datasets used to evaluate rudimentary models include the Iris dataset with 150 samples [17] and the Boston House Prices dataset with 506 instances [18]. Even within the realm of medical research, valuable datasets such as the breast cancer dataset encompass only around 300 patients [19]. Smaller datasets pose challenges for DGM training, including overfitting and difficulty in assessing the quality of generated data [20].
@@
A critical challenge associated with using DGMs for tabular data generation lies in ensuring the quality and effectiveness of synthetic data. While standardized metrics exist for image [21, 22] and text data [23, 24], measuring the quality of synthetic tabular data presents unique challenges. Studies employ various metrics, including pairwise correlation difference, support coverage, likelihood fitness, and other statistics as described in [25]. However, a consistent method for holistic evaluation is lacking. Divergences, which quantify the discrepancy between probability distributions, offer a promising avenue for validation [20]. They can capture the overall differences between real and synthetic data by considering the joint distribution of all attributes. However, modeling joint distributions presents a trade-off between computational cost and accuracy. Large datasets, especially those with high dimensionality, require significant computational resources. With sufficient resources, accurate results can be achieved. In contrast, smaller datasets, common in real-world applications, present a challenge to accuracy. Limited data may hinder the capture of complex variable relationships, leading to models with poor generalization to unseen data. Consequently, even computationally efficient methods for joint distribution modeling can yield inaccurate results in small data settings.
@@
The limitations of current validation techniques further compound the inherent limitations of small datasets. These techniques often focus on comparing synthetic data with real data used for training, failing to account for the limited scope of information on which the DGM was trained. This can lead to a false sense of security, where the synthetic data appear similar to the training data but may not generalize well. Deep learning models with high parameter counts are susceptible to overfitting too, especially on small datasets. Additionally, small datasets might not capture the full spectrum of real-world variations. Consequently, the synthetic data generated may not be representative of the underlying distribution, affecting its effectiveness for different tasks. Furthermore, smaller datasets are more prone to the influence of noise (random errors) and bias (systematic skews). These mislead DGMs into learning incorrect patterns, ultimately resulting in unrealistic synthetic data.
@@
This work addresses the critical challenge of generating reliable synthetic tabular data from limited datasets, a prevalent scenario in many real-world applications. Traditional divergence metrics often struggle in these situations, which can lead to inaccurate assessments of the quality of synthetic data [20]. We propose a novel methodology specifically designed to address this issue by introducing a framework that leverages inductive biases to improve the performance of DGMs in small dataset environments. Inductive biases [26] are inherent preferences or assumptions built into a learning model. These biases can guide the learning process and improve model performance, particularly when data is scarce. Traditionally, inductive biases are introduced through domain knowledge or specific architectural choices. This work proposes an alternative approach that leverages the variability that is often found in the DGM training process to generate inductive biases through different learning techniques. As an example, we use VAEs, as they exhibit inherent variability between training seeds, allowing them to capture various aspects of data during training, but we note that our approach could be used with other DGMs. We exploit this variability by employing various transfer learning and meta-learning techniques to generate the inductive bias, ultimately leading to improved synthetic data generation. Our key contribution is threefold:
@@
  • We propose four different techniques (pre-training, model averaging, Model-Agnostic Meta-Learning (MAML), and Domain-Randomized Search (DRS)) to generate the inductive bias. We demonstrate the efficacy of our proposal using a common DGM as the VAE, and for the pre-training process, we also assess the performance with a CTGAN architecture, another common DGM.
  • We demonstrate the effectiveness of our proposed methodology through extensive experiments on benchmark datasets. We provide a comprehensive analysis of the results using established divergence metrics such as Kullback-Leibler (KL) and Jensen-Shannon (JS) divergences.
This contribution offers a significant advancement in the field of synthetic data generation. By enabling improved data quality in resource-constrained settings, our approach has the potential to broaden the applicability of synthetic tabular data across various disciplines.
@@ 2 Methodology: Generating Artificial Inductive Bias
Let us assume that we have a tabular dataset composed of $N$ entries $\{x_{r}^{i}\}_{i=1}^{N}$. $N$ represents the number of samples available and each entry $x_{r}^{i}$ has a dimensionality of $C$ features. In other words, $C$ represents the number of attributes associated with each data point. Let us also define a DGM as a high-dimensional probability distribution $p_{\theta}$, where $\theta$ represents the learnable parameters of the model. The objective of the DGM is to learn a representation, $p_{\theta}$, that closely approximates the true underlying data distribution, denoted by $p(x_{r})$. Once trained, the DGM can generate new synthetic generated samples $x_{g}$, by drawing from its learned distribution:
@@
In the prevalent big data setting, characterized by a large number of training samples ($N\gg C$), DGMs with sufficient complexity can effectively capture the underlying data distribution $p(x_{r})$. This is evidenced by the impressive results achieved in recent research, where high-dimensional synthetic samples are generated using vast amounts of training data [7, 1, 2, 3]. However, for scenarios with limited training data, which is common in tabular domains, DGMs struggle to accurately represent the complex inter-feature relationships. Consequently, the synthetic samples generated $x_{g}$ deviate significantly from the true data distribution $p(x_{r})$, leading to high KL and JS divergences between real and synthetic data.
@@
To address this challenge, we propose an approach that uses artificially generated inductive biases. Figure 1 illustrates the overall architecture. In the standard big data setting, a DGM $p_{\theta}$ is directly trained using real data $x_{r}$ generating high-quality synthetic data $x_{g}\sim p_{\theta}$. However, when the number of real samples $N$ is limited, the quality of the generated data $x_{g}$ deteriorates. To mitigate this issue, we introduce an artificial inductive bias generator. This module takes the initial synthetic data $x_{g}$ as input and outputs an initial set of weights $\theta_{0}$. These weights are then used as the inductive bias to train a second DGM $p_{\hat{\theta}}$ using real data $x_{r}$. This second DGM generates a new set of synthetic samples, $\hat{x}_{g}$. Notably, the only distinction between $p_{\theta}$ and $p_{\hat{\theta}}$ lies in the initial weights: $p_{\hat{\theta}}$ leverages the inductive bias encoded in $\theta_{0}$ to potentially achieve faster convergence to a distribution that better resembles $p(x_{r})$, while $p_{\theta}$ begins training with random weights. As our simulations will demonstrate, this seemingly minor difference translates into significant improvements in the quality of the generated synthetic data.
@@
Figure 1: Block diagram for the proposed architecture. In a standard big data setting, the first generative model $p_{\theta}$ generates good enough samples. However, in cases where large amounts of data are not available ($N$ is limited), we propose to use the data generated by the first DGM as input to an artificial inductive bias generator, which in return provides a set of initial weights $\theta_{0}$ for a DGM that contains the artificially generated inductive bias. This initial weight is then used to train a second DGM $p_{\hat{\theta}}$ using real data $x_{r}$, generating a second set of synthetic data $\hat{x}_{g}$, which is of higher quality than the synthetic data $x_{g}$.
@@
The proposed approach hinges on two key concepts: the importance of inductive biases and the feasibility of their artificial generation. The importance of inductive biases in supervised learning is well established. The no-free-lunch theorems state that a universally optimal learner does not exist. Consequently, specific learning biases can lead to substantial performance gains for particular problem domains (see [27] and the references therein). Convolutional Neural Networks (CNNs) exemplify this principle. Their inherent inductive bias, the fact that the image information possesses spatial correlation, makes them the preferred architecture for image processing tasks. Similarly, as highlighted in [26], the use of inductive biases is a cornerstone of Deep Learning’s success. In scenarios with limited training data, regularizers are commonly employed as inductive biases to prevent overfitting. This underscores the dual role of inductive biases: not only do they contribute to Deep Learning’s effectiveness, but they are also crucial in preventing overfitting. However, effective use of inductive biases is often contingent on having specific knowledge about the problem at hand. In the aforementioned example of CNNs, we inherently understand the existence of spatial correlation in images. However, in tabular data, this domain-specific knowledge is often scarce. To address this challenge, recent efforts have focused on designing large models trained on artificially generated data as inductive biases. The underlying hope is that the actual problem to be solved exhibits similarities to those encountered during training of the large model (e.g., [28] and [29]).
@@
Therefore, our proposed approach departs from existing methods for incorporating inductive biases in synthetic tabular data generation. Unlike the brute-force approach employed in [29], we use data generated by a potentially low-quality DGM $p_{\theta}$. This strategy aims to obtain an initial set of weights $\theta_{0}$, which act as an inductive bias. Ideally, these weights should guide the model to a region of the parameter space that facilitates convergence towards a high-quality solution. IIn particular, we assess our ideas using a state-of-the-art VAE architecture for the DGM, although we hypothesize that similar results could be achieved with other DGM architectures. Due to its demonstrated superiority against other leading models, we will use the architecture proposed in [15]. VAEs are known to be sensitive to the initial random conditions (seeds) used during training. This dependence on seeds requires training with multiple seeds and selecting the one(s) that exhibit the best performance based on a chosen metric, such as minimum validation loss. The remaining runs, often discarded, may still contain valuable problem-specific information despite not achieving optimal solutions using traditional metrics. Our key idea lies in exploiting these potentially informative data from discarded VAE runs to create an artificial inductive bias for the final DGM trained with real data.
@@
The following subsections explore two distinct paradigms for generating the initial set of weights, $\theta_{0}$: transfer learning and meta-learning. Transfer learning techniques encompass pre-training and model averaging, while meta-learning techniques include MAML and DRS. Pre-training offers a versatile approach applicable to any DGM architecture, regardless of inherent characteristics. In contrast, model averaging and meta-learning techniques are particularly well suited for VAEs trained with multiple seeds due to their inherent variability in learned representations. Consequently, we will evaluate the latter two techniques within the chosen VAE architecture. Additionally, to assess the efficacy of pre-training across different DGM architectures, we will compare its performance on the CTGAN.

### v2

@@ Abstract
While synthetic tabular data generation using Deep Generative Models (DGMs) offers a compelling solution to data scarcity and privacy concerns, their effectiveness relies on the availability of substantial training data, often lacking in real-world scenarios. To overcome this limitation, we propose a novel methodology that explicitly integrates artificial inductive biases into the generative process to improve data quality in low-data regimes. Our framework leverages transfer learning and meta-learning techniques to construct and inject informative inductive biases into DGMs. We evaluate four approaches (pre-training, model averaging, Model-Agnostic Meta-Learning (MAML), and Domain Randomized Search (DRS)) and analyze their impact on the quality of the generated text. Experimental results show that incorporating inductive bias substantially improves performance, with transfer learning methods outperforming meta-learning, achieving up to 60% gains in Jensen-Shannon divergence. The methodology is model-agnostic and especially relevant in domains such as healthcare and finance, where high-quality synthetic data are essential, and data availability is often limited.
@@ 1 Introduction
Recent advances in Deep Learning (DL) have led to the development of powerful Deep Generative Models (DGMs). These models excel at learning and representing complex, high-dimensional distributions, allowing them to sample new data points realistically. This capability has driven remarkable progress in various domains, including image generation (elasri2022image, ), text generation (zhang2023survey, ), and video generation (selva2023video, ).
Tabular data has gained increasing interest within DGMs due to their structured format, making them fundamental for information storage and analysis across various fields. Researchers are actively investigating techniques for generating realistic and informative tabular data, as evidenced by the growing work in this area (borisov2022deep, ; fonseca2023tabular, ). Since the introduction of Generative Adversarial Networks (GANs) (Goodfellow2014GenerativeAN, ), a wide range of GAN-based models have been proposed for generating synthetic tabular data. Among them, CTGAN (xu2019modeling, ) is one of the most prominent and widely adopted models, as it introduces conditional generation mechanisms that allow it to handle the multimodal, imbalanced, and mixed-type nature of tabular data. Building on CTGAN, CTAB-GAN (zhao2021ctab, ) further improves the generation of both categorical and continuous variables with complex distributions. GANBLR (zhang2021ganblr, ) proposes an interpretable GAN-based generator tailored to categorical data, achieving strong performance in both statistical similarity and downstream utility. Another notable architecture is TableGAN (park2018data, ), which enhances vanilla GANs by incorporating classification and information loss objectives to improve the semantic and statistical alignment between real and synthetic samples. However, TableGAN is less effective at handling categorical imbalances and lacks the conditional generation capabilities of CTGAN-based models. In addition, the ACGAN (odena2017conditional, ) extends the discriminator to include an auxiliary classification task, enabling class-aware generation and improving semantic consistency. While originally designed for in-distribution generation, ACGAN has served as the foundation for more recent frameworks that generalize beyond the training distribution. For instance, (jian2024open, ) proposes an open-set domain generalization framework that extends ACGAN with dual-level weighting mechanisms to synthesize fault diagnosis data under unseen operating conditions. Similarly, (jian2024gradient, ) introduces a gradient-based meta-learning strategy that augments the source domain to improve single-domain generalization. These works operate in the out-of-distribution (OOD) setting, where the objective is to generate or adapt to novel data distributions not present during training. In contrast, our work focuses on generating high-quality in-distribution data under data scarcity conditions. Rather than expanding beyond the original data distribution, we aim to faithfully reproduce it, particularly in critical domains such as healthcare and finance, where maintaining the statistical integrity of the data is essential.
Beyond GANs, Variational Autoencoders (VAEs) (Kingma2013AutoEncodingVB, ) have emerged as a robust alternative for generating tabular data. Models such as TVAE (xu2019modeling, ) are specifically designed to manage mixed data types, while HI-VAE (nazabal2020handling, ) focuses on handling heterogeneous and incomplete datasets. VAEM (ma2020vaem, ) extends this line by integrating imputation and generative modeling. More recently, the VAE-BGM (apellaniz2024improved, ) has integrated a Bayesian Gaussian Mixture (BGM) in the VAE architecture, replacing the standard isotropic Gaussian assumption in the latent space (Cremer2018InferenceSI, ). This modification enables the model to sample from a more expressive latent space, generating higher-quality data, particularly in situations with limited availability. Additionally, emerging approaches based on diffusion models (ho2020denoising, ) and transformer architectures have shown promise in generating synthetic tabular data. Models such as STaSy (kim2022stasy, ) and TabSyn (zhang2023mixed, ) apply score-based generative modeling with latent encoding strategies. In contrast, TabDDPM (kotelnikov2023tabddpm, ) performs diffusion in both Gaussian and multinomial spaces to effectively model continuous and categorical features. These models benefit from improved training stability and are particularly effective in capturing complex inter-feature dependencies.
@@
The importance of sufficient data for DGM training cannot be overstated. Studies using popular DGMs, such as CTGAN, utilize datasets ranging from 10,00010,00010 , 000 to 481,000481,000481 , 000 training instances. This contrasts starkly with the practices observed in numerous domains. For example, well-established datasets used to evaluate rudimentary models include the Iris dataset, with 150 samples (misc_iris_53, ), and the Boston House Prices dataset, with 506506506 instances (HARRISON197881, ). Even within the realm of medical research, valuable datasets such as the breast cancer dataset encompass only around 300 patients (misc_breast_cancer_14, ). Smaller datasets pose challenges for DGM training, including overfitting and difficulty assessing the quality of generated data (apellaniz2024synthetic, ).
@@
A critical challenge associated with using DGMs for generating tabular data lies in ensuring the quality and effectiveness of the synthetic data. While standardized metrics exist for image (heusel2018gans, ; salimans2016improved, ) and text data (zhang2020bertscore, ; Papineni2002BleuAM, ), measuring the quality of synthetic tabular data presents unique challenges. Studies employ various metrics, including pairwise correlation difference, support coverage, likelihood fitness, and other statistics described in (vicomtech, ). However, a consistent method for holistic evaluation is lacking. Divergences, which quantify the discrepancy between probability distributions, offer a promising avenue for validation (apellaniz2024synthetic, ). They can capture the overall differences between real and synthetic data by considering the joint distribution of all attributes. However, modeling joint distributions presents a trade-off between computational cost and accuracy. Large datasets, especially those with high dimensionality, require significant computational resources. With sufficient resources, accurate results can be achieved. In contrast, smaller datasets, common in real-world applications, present a challenge to accuracy. Limited data may hinder the capture of complex variable relationships, leading to models with poor generalization to unseen data. Consequently, even computationally efficient methods for joint distribution modeling can yield inaccurate results in small data settings.
@@
The limitations of current validation techniques further compound the inherent limitations of small datasets. These techniques often focus on comparing synthetic data with real data used for training, failing to account for the limited scope of information on which the DGM was trained. This can lead to a false sense of security, where the synthetic data appear similar to the training data but may not generalize well. DL models with high parameter counts are susceptible to overfitting, too, especially on small datasets. Additionally, small datasets might not capture the full spectrum of real-world variations. Consequently, the synthetic data generated may not accurately represent the underlying distribution, which can affect its effectiveness for various tasks. Furthermore, smaller datasets are more prone to the influence of noise (random errors) and bias (systematic skews). These mislead DGMs into learning incorrect patterns, ultimately resulting in the generation of unrealistic synthetic data.
@@
This work addresses the critical challenge of generating reliable synthetic tabular data from limited datasets, a prevalent scenario in many real-world applications. Traditional divergence metrics often struggle in these situations, leading to inaccurate assessments of the quality of synthetic data (apellaniz2024synthetic, ). We propose a novel methodology specifically designed to address this issue by introducing a framework that leverages inductive biases to improve the performance of DGMs in small dataset environments. Inductive biases (goyal2022inductive, ) are inherent preferences or assumptions built into a learning model. These biases can guide learning and improve model performance, particularly when data are scarce. Traditionally, inductive biases are introduced through domain knowledge or specific architectural choices. This work proposes an alternative approach that leverages the variability often found in the DGM training process to generate inductive biases through different learning techniques. For example, we use VAEs, as they exhibit inherent variability between training seeds, allowing them to capture various aspects of data during training. Still, we note that our approach could be applied to other DGMs. We exploit this variability by employing various transfer learning and meta-learning techniques to generate the inductive bias, ultimately leading to improved synthetic data generation. Our key contribution is threefold:
@@
  • We propose four different techniques (pre-training, model averaging, Model-Agnostic Meta-Learning (MAML), and Domain Randomized Search (DRS)) to generate the inductive bias. We demonstrate the efficacy of our proposal using a common DGM as the VAE. We also assess the performance with a CTGAN architecture, another common DGM for the pre-training process.
  • We conduct extensive experiments on benchmark datasets and evaluate the quality of the generated data using both divergence-based metrics, such as the Kullback-Leibler (KL) and Jensen-Shannon (JS) divergences, and a utility-based validation protocol. This latter validation involves training a downstream model on synthetic data and testing it on real data, thereby measuring the practical utility of the generated samples for predictive tasks.
@@ 2 Methodology: Generating Artificial Inductive Bias
Assuming a tabular dataset composed of $N$ entries $\{x_{r}^{i}\}_{i=1}^{N}$, where $N$ represents the number of samples available and each entry $x_{r}^{i}$ has a dimensionality of $C$ features. In other words, $C$ represents the number of attributes associated with each data point. Let us also define a DGM as a high-dimensional probability distribution $p_{\theta}$, where $\theta$ represents the learnable parameters of the model. The objective of the DGM is to learn a representation, $p_{\theta}$, that closely approximates the true underlying data distribution, denoted by $p(x_{r})$. Once trained, the DGM can generate new synthetic samples $x_{g}$ by drawing from its learned distribution:
@@
In the prevalent big data setting, characterized by a large number of training samples ($N\gg C$), DGMs with sufficient complexity can effectively capture the underlying data distribution $p(x_{r})$. This is evidenced by the impressive results achieved in recent research, where high-dimensional synthetic samples are generated using vast amounts of training data (xu2019modeling, ; elasri2022image, ; zhang2023survey, ; selva2023video, ). However, for scenarios with limited training data, which is common in tabular domains, DGMs struggle to accurately represent the complex inter-feature relationships. Consequently, the synthetic samples generated $x_{g}$ deviate significantly from the true data distribution $p(x_{r})$, leading to high KL and JS divergences between real and synthetic data.
@@
We propose an approach that uses artificially generated inductive biases to address this challenge. Fig. 1 illustrates the overall architecture. In the standard big data setting, a DGM $p_{\theta}$ is directly trained using real data $x_{r}$, generating high-quality synthetic data $x_{g}\sim p_{\theta}$. However, when the number of real samples $N$ is limited, the quality of the generated data $x_{g}$ deteriorates. To mitigate this issue, we introduce an artificial inductive bias generator. This module inputs the initial synthetic data $x_{g}$ and outputs an initial set of weights $\theta_{0}$. These weights are then used as the inductive bias to train a second DGM $p_{\hat{\theta}}$ using real data $x_{r}$. This second DGM generates a new set of synthetic samples, $\hat{x}_{g}$. Notably, the only distinction between $p_{\theta}$ and $p_{\hat{\theta}}$ lies in the initial weights: $p_{\hat{\theta}}$ leverages the inductive bias encoded in $\theta_{0}$ to potentially achieve faster convergence to a distribution that better resembles $p(x_{r})$. At the same time, $p_{\theta}$ begins training with random weights. As our simulations will demonstrate, this seemingly minor difference translates into significant improvements in the quality of the generated synthetic data.
@@
Figure 1: Block diagram of the proposed architecture. In standard big-data settings, the first generative model $p_{\theta}$ generates samples. However, when data are scarce ($N$ is limited), we use its output to create an artificial inductive bias, providing initial weights $\theta_{0}$ for a second DGM $p_{\hat{\theta}}$. This second model is then trained on real data $x_{r}$, producing higher-quality synthetic data $\hat{x}_{g}$ than the initial synthetic samples $x_{g}$.
@@
The proposed approach hinges on two key concepts: the importance of inductive biases and the feasibility of generating them artificially. The importance of inductive biases in supervised learning is well established. The no-free-lunch theorems state that a universally optimal learner does not exist. Consequently, specific learning biases can lead to substantial performance gains for particular problem domains (see goldblum2023no and the references therein). Convolutional Neural Networks (CNNs) exemplify this principle. Their inherent inductive bias, which is the fact that the image information possesses spatial correlation, makes them the preferred architecture for image processing tasks. Similarly, as highlighted in (goyal2022inductive, ), using inductive biases is a cornerstone of DL’s success. In scenarios with limited training data, regularizers are commonly employed as inductive biases to prevent overfitting. This underscores the dual role of inductive biases: not only do they contribute to DL’s effectiveness, but they are also crucial in avoiding overfitting. However, effective use of inductive biases is often contingent on having specific knowledge about the problem. In the aforementioned example of CNNs, we inherently understand spatial correlation in images. However, in tabular data, this domain-specific knowledge is often scarce. Recent efforts have focused on designing large models trained on artificially generated data as inductive biases to address this challenge. The underlying hope is that the actual problem to be solved exhibits similarities to those encountered during training of the large model (e.g., muller2021transformers and hollmann2022tabpfn ).
Therefore, our proposed approach departs from existing methods for incorporating inductive biases in synthetic tabular data generation. Unlike the brute-force approach employed in hollmann2022tabpfn , we use data generated by a potentially low-quality DGM $p_{\theta}$. This strategy aims to obtain an initial set of weights $\theta_{0}$, which act as an inductive bias. Ideally, these weights should guide the model to a region of the parameter space that facilitates convergence towards a high-quality solution.
@@
In particular, we assess our ideas using a state-of-the-art VAE architecture for the DGM, although we hypothesize that similar results could be achieved with other DGM architectures. Due to its demonstrated superiority against other leading models, we will use the architecture proposed in apellaniz2024improved . VAEs are known to be sensitive to the initial random conditions (seeds) used during training. This dependence on seeds requires training with multiple seeds and selecting the one (or more) that exhibit the best performance based on a chosen metric, such as the minimum validation loss. The remaining runs, often discarded, may still contain valuable problem-specific information despite not achieving optimal solutions using traditional metrics. Our key idea lies in exploiting the potentially informative data from discarded VAE runs to create an artificial inductive bias for the final DGM, which is trained with real data. In addition to its demonstrated empirical performance, this VAE-based model incorporates architectural elements that inherently discourage overfitting and sample memorization, an important consideration in data-scarce regimes. Specifically, the model introduces a two-stage sampling process: first, latent representations are sampled from the encoder’s posterior distribution; then, instead of decoding directly from these individual latent points, a BGM is fitted over the latent means, from which a second sample is drawn before decoding. This structure increases diversity in the generated samples and reduces the likelihood of memorizing specific training records. Moreover, the VAE formulation enforces a KL divergence regularization between the approximate posterior and the prior, further penalizing overly confident or localized representations that could lead to overfitting. These design choices serve as strong inductive biases that favor generalization over replication, complementing our overall strategy.
@@
The following subsections explore two distinct paradigms for generating the initial set of weights, $\theta_{0}$: transfer learning and meta-learning. Transfer learning techniques encompass pre-training and model averaging, while meta-learning techniques include MAML and DRS. Pre-training offers a versatile approach applicable to any DGM architecture, regardless of its inherent characteristics. In contrast, model averaging and meta-learning techniques are particularly well-suited for VAEs trained with multiple seeds due to their intrinsic variability in learned representations. Consequently, we will evaluate the latter two methods within the chosen VAE architecture. Additionally, to assess the efficacy of pre-training across different DGM architectures, we will compare its performance on the CTGAN.